OpenStudy (anonymous):

-3(2-x) +1 = 4 - (7x-2)

OpenStudy (cggurumanjunath):

-3(2-X)=-3*2+(-3*-X)=-6+3X

OpenStudy (cggurumanjunath):

IS IT CLEAR @Jennalylnn ?

OpenStudy (cggurumanjunath):